### 05 Mar 2026 v2

# Predictive Coherence and the Moment Hierarchy: Martingale Posteriors for Exchangeable Bernoulli Sequences

Abstract
For an exchangeable Bernoulli sequence with de Finetti mixing measure $\Pi$, the $k$-step predictive probability $\mathbb{P}(X_{n+1}=\cdots=X_{n+k}=0\mid\mathcal{F}_{n})$ equals the posterior expectation $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$. By binomial expansion, this depends on all posterior moments up to order $k$. We show that the first moment alone is not sufficient to uniquely identify these quantities: for $k\geq 2$, the mapping from posterior mean to $k$-step predictive is set-valued. The martingale posterior framework of Fong, Holmes, and Walker (which constrains only the first conditional moment of the terminal value) does not, in general, uniquely identify multi-step predictive distributions. Under any strictly proper scoring rule, the plug-in predictive is strictly dominated by the Bayes predictive whenever the posterior is non-degenerate. A closure theorem establishes that a martingale posterior determines all $k$-step predictives if and only if the conditional law of the terminal value is uniquely specified. Hill’s $A_{(n)}$ rule under the Jeffreys $\mathrm{Beta}(\tfrac{1}{2},\tfrac{1}{2})$ prior is a positive example. The discrepancy is $O(\mathrm{Var}(\theta\mid\mathcal{F}_{n}))$ and vanishes as the posterior concentrates. These results clarify the structural requirements for predictive completeness under exchangeability.
1 Introduction
Let $(X_{i})_{i\geq 1}$ be an exchangeable $\{0,1\}$-valued sequence. By de Finetti’s theorem, there exists a unique mixing measure $\Pi$ on $[0,1]$ such that, conditionally on $\theta\sim\Pi$, the $X_{i}$ are i.i.d. $\mathrm{Bern}(\theta)$. The posterior $\Pi(\cdot\mid\mathcal{F}_{n})$ determines all predictive quantities through integration.
Fong, Holmes, and Walker (2023) proposed replacing the prior-times-likelihood mechanism with a single coherence condition: the sequence $(\theta_{n})_{n\geq 0}$ must satisfy
$\mathbb{E}[\theta_{n}\mid\mathcal{F}_{n-1}]=\theta_{n-1}\quad\text{a.s.},\qquad n\geq 1,$ (1)
where $\mathcal{F}_{n}=\sigma(X_{1:n})$ and $\mathcal{F}_{0}$ is trivial. By the martingale convergence theorem, $\theta_{n}\to\theta_{\infty}$ a.s., and the law of the directing parameter $\theta_{\infty}$ is a posterior. Condition (1) constrains only the first conditional moment of $\theta_{\infty}$ given $\mathcal{F}_{n}$.
We study the extent to which first-moment coherence determines multi-step predictive quantities. The Bernoulli case is the natural setting: it isolates the structural phenomenon in the simplest exchangeable model.
Theorem (Main results, informal).
The martingale condition identifies the posterior mean; we characterize what additional structure is required for predictive completeness. Specifically: for $k\geq 2$, the mapping $m_{n}\mapsto\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ is set-valued, so first-moment information does not uniquely identify the $k$-step predictive. Under any strictly proper scoring rule, the plug-in rule is strictly dominated by the Bayes predictive whenever the posterior is non-degenerate. A martingale posterior is predictively complete if and only if the conditional law of $\theta_{\infty}$ is uniquely specified.
Goldstein’s (1975a) conditional-prevision programme organises beliefs at the moment level: the analyst commits to selected expectation functionals rather than a full posterior law. Cover’s (1985) pathwise programme accumulates log-evidence at the KL rate, supplying the e-process structure for anytime-valid inference. The martingale posterior of Fong, Holmes, and Walker (2023) imposes a first-moment coherence condition analogous to Goldstein’s prevision commitment; its almost-sure convergence connects it to Cover’s framework. Table 1 summarises these relationships.
Table 1: Four frameworks for posterior belief about an exchangeable sequence, ordered by strength of commitment. Commitment is what the analyst is required to specify; posterior object is what is thereby determined; multi-step uniqueness asks whether block-event predictives are uniquely fixed for every horizon. The non-uniqueness result for the final row is Theorem 6.3.
Framework Commitment Posterior object Multi-step uniqueness
de Finetti coherence Full prior on directing measure Full posterior law Yes (de Finetti’s theorem)
Classical Bayesian updating Likelihood and prior Full posterior law Yes (Bayes’ rule)
Goldstein conditional previsions Chosen expectation functionals (conditional previsions) A finite set of posterior moments Partial: each committed prevision fixes one predictive horizon
Martingale posteriors (this paper) First conditional moment only Terminal-value law No: underdetermined for every horizon beyond one
The first two frameworks inherit full predictive uniqueness from their commitment to a complete posterior law. The Goldstein row makes explicit that committing to finitely many expectation functionals determines only the corresponding predictive horizons. The final row identifies the case studied here: Theorems 6.3 and 10.3 characterise the resulting gap and its closure condition.
De Finetti’s theorem and Sanov’s theorem are dual descriptions of the same sample path. De Finetti provides a directing measure $\mu$ as the almost-sure limit of the empirical measure $L_{n}$; Sanov quantifies the cost of deviating from that limit:
$\mathbb{P}(L_{n}\in A)\;\asymp\;\exp\!\Bigl(-n\inf_{\mu\in A}D_{\mathrm{KL}}(\mu\|\mu_{0})\Bigr).$
Reversing the KL argument (a standard reweighting identity; see e.g. Cover and Thomas 2006, Ch. 12) gives the inverse Sanov identity: posterior mass near $\mu$ concentrates at rate $\exp(-n\,D_{\mathrm{KL}}(L_{n}\|\mu))$, recovering the Bernstein–von Mises theorem as the moderate-deviations limit. The martingale posterior of Fong, Holmes, and Walker (2023) sits at the intersection of both: $(\Pi_{n})$ is a measure-valued martingale whose almost-sure limit $\Pi_{\infty}=\delta_{\mu}$ recovers the de Finetti directing measure, while its contraction rate is governed by $D_{\mathrm{KL}}(\cdot\|\mu_{0})$. Section 2 develops this connection; Section 4 shows that first-moment information does not determine the higher-order structure of $\Pi_{\infty}$ (Theorem 6.3).
Goldstein (1975a) and Goldstein and Wooff (1998) develop the Bayes linear programme in which conditional previsions are the primary inferential object. In a companion paper, Goldstein (1975b) showed that when the posterior mean is a linear function of the sample, a recursive relation uniquely determines all moments of the prior distribution—provided the likelihood moments are known. The present paper establishes that without this likelihood structure, first-moment constraints alone do not identify block-event predictives (Theorem 6.3). Diaconis and Ylvisaker (1979) characterized conjugate priors by linearity of the posterior mean; Condition (1) is the sequential analogue. Polson and Zantedeschi (2025) construct predictive distributions from finite moment constraints; Datta and Polson (2025) show polynomial log-marginals of degree $\geq 3$ cannot correspond to any valid prior. Section 4 identifies which multi-step predictives remain underdetermined under first-moment constraints (Theorem 6.3).
The main results are the moment hierarchy and insufficiency theorem (Theorems 4.1 and 6.3), the plug-in dominance result (Proposition 7.3), and a closure theorem characterizing predictive completeness (Theorem 10.3). Hill’s $A_{(n)}$ rule under the Jeffreys $\mathrm{Beta}(\tfrac{1}{2},\tfrac{1}{2})$ prior serves as a positive example (Section 8). A companion paper (Polson and Zantedeschi, 2026) develops a broader taxonomy of admissibility geometries (Bayes risk, e-process validity, conformal coverage, and Cesàro approachability); the present paper isolates the moment-hierarchy obstruction behind mean-only coherence.
The remainder of the paper is organised as follows. Section 2 develops the sample-path framework and Section 3 sets up the Bernoulli model. Sections 4 through 7 establish the moment hierarchy, insufficiency, and plug-in dominance. Section 8 provides a positive example via Hill’s $A_{(n)}$ rule. Sections 9 and 10 give asymptotic reconciliation and the closure theorem. Section 11 treats optimal stopping, and Section 12 concludes.
2 Sanov, de Finetti, and the sample-path picture
This section outlines a sample-path viewpoint connecting KL rates, exchangeability, and posterior concentration; it provides context for the Bernoulli results and is not required for the main theorems. The remainder of the paper makes the Bernoulli case concrete.
Given i.i.d. observations $X_{1},\ldots,X_{n}$ with law $\mu_{0}$, the empirical measure $L_{n}=n^{-1}\sum_{i=1}^{n}\delta_{X_{i}}$ satisfies a large-deviation principle (LDP) with speed $n$ and rate function
$I(\mu)=D_{\mathrm{KL}}(\mu\|\mu_{0})=\int\log\frac{d\mu}{d\mu_{0}}\,\mathrm{d}\mu.$ (2)
The probability that $L_{n}$ lies in a measurable set $A$ decays as
$\mathbb{P}(L_{n}\in A)\asymp\exp\!\bigl(-n\inf_{\mu\in A}D_{\mathrm{KL}}(\mu\|\mu_{0})\bigr)$
(Sanov, 1957). The KL divergence is the canonical cost of deviation from the truth.
For an exchangeable sequence $(X_{i})_{i\geq 1}$, de Finetti’s theorem (Theorem 3.2 below) provides a unique prior $\Pi$ on the space of data-generating laws such that, conditionally on $\mu\sim\Pi$, the observations are i.i.d. with law $\mu$. By the strong law, $L_{n}\to\mu$ a.s., so the law of the limit is $\mathcal{L}(L_{\infty})=\Pi$. Each infinite sample path reveals exactly one draw from $\Pi$; the prior itself is not observable from a single path.
There is a duality between the forward and backward directions of the rate function. In the forward direction (Sanov), the cost of observing $L_{n}\approx\nu$ when the truth is $\mu_{0}$ is $\approx\exp(-n\,D_{\mathrm{KL}}(\nu\|\mu_{0}))$. In the inverse direction, fixing the data $L_{n}$, posterior mass near $\mu$ concentrates at rate $\exp\{-n\,D_{\mathrm{KL}}(L_{n}\|\mu)\}$. The KL arguments are reversed. The posterior $\Pi_{n}$ is the conditional law of $\mu$ given $L_{n}$, interpolating $\Pi_{0}=\Pi\;\longrightarrow\;\Pi_{\infty}=\delta_{\mu}$. At moderate deviations (scale $a_{n}/\sqrt{n}$ with $a_{n}\to\infty$, $a_{n}=o(\sqrt{n})$), the Sanov rate is approximated by the quadratic
$D_{\mathrm{KL}}(\mu\|\mu_{0})\approx\tfrac{1}{2}(\theta-\theta_{0})^{T}\mathcal{I}(\theta_{0})(\theta-\theta_{0}),$ (3)
where $\mathcal{I}(\theta_{0})$ is the Fisher information matrix. Eichelsbacher and Ganesh (2002) show the posterior satisfies a moderate-deviation principle (MDP) with this quadratic rate, sharpening the Bernstein–von Mises theorem to intermediate scales. For an extension of Sanov’s theorem to Gibbs conditioning and tilted empirical measures, see Léonard and Najim (2002).
An e-variable is a nonnegative random variable $E$ with $\mathbb{E}_{H_{0}}[E]\leq 1$. An e-process $(E_{n})$ is a test martingale under $H_{0}$; Ville’s inequality gives type-I error control at any stopping time (Ville, 1939; Grünwald et al., 2024). The log of the growth-rate-optimal e-process satisfies
$\frac{1}{n}\log E_{n}\;\longrightarrow\;D_{\mathrm{KL}}(\mu\|\mu_{0})\quad\text{a.s.}$ (4)
When $\mu=\mu_{0}$, $E_{n}$ is a martingale and does not drift; when $\mu\neq\mu_{0}$, $E_{n}\to\infty$ at exponential rate $D_{\mathrm{KL}}(\mu\|\mu_{0})>0$. The alternative is discovered from the sample path as $L_{n}$ converges, not specified in advance.
The Bayes factor $E_{n}=\int p(X_{1:n}\mid\mu)\Pi(d\mu)\big/p(X_{1:n}\mid\mu_{0})$ is automatically a test martingale under $H_{0}$. The prior $\Pi$ affects power but not validity.
Table 2 collects the four frameworks. Each describes the information in a single sample path about the directing measure $\mu$.
Table 2: Four descriptions of sample-path information.
Framework Primary object Path convergence Rate
De Finetti Directing measure $\mu\sim\Pi$ $L_{n}\to\mu$ a.s. Strong law
Sanov/LDP Empirical measure $L_{n}$ $\mathbb{P}(L_{n}\in A)\asymp e^{-nI}$ $D_{\mathrm{KL}}(\cdot\|\mu_{0})$
E-process Test martingale $E_{n}$ $n^{-1}\log E_{n}\to D_{\mathrm{KL}}$ Linear in $n$
Mart. posterior Measure-valued mg. $\Pi_{n}$ $\Pi_{n}\to\delta_{\mu}$ a.s. Posterior contraction
The martingale posterior of Fong, Holmes, and Walker (2023) sits in the bottom row: it constrains the first moment of $\Pi_{n}$ but not the law. The insufficiency results of Section 6 quantify this gap; the closure theorem (Theorem 10.3) identifies the exact boundary.
The following theorem makes the Sanov–de Finetti duality concrete.
Theorem 2.1 (Sanov–Moment Bridge).
Let $(X_{i})_{i\geq 1}$ be exchangeable Bernoulli with prior $\Pi_{0}$ and empirical mean $L_{n}=S_{n}/n$.
  1. (i)
    Posterior as Sanov exponential. The Bernoulli likelihood satisfies $\theta^{S_{n}}(1-\theta)^{n-S_{n}}=\exp\{-n\,D_{\mathrm{KL}}(L_{n}\|\theta)+nH(L_{n})\}$, where $H(L_{n})$ is the binary entropy at $L_{n}$. Absorbing the $\mathcal{F}_{n}$-measurable factor $\exp\{nH(L_{n})\}$ into the normalising constant, the posterior is
    $\Pi_{n}(d\theta)\;\propto\;\exp\!\bigl\{-n\,D_{\mathrm{KL}}(L_{n}\|\theta)\bigr\}\,\Pi_{0}(d\theta).$ (5)
    The posterior is shaped entirely by the Sanov rate function $\theta\mapsto D_{\mathrm{KL}}(L_{n}\|\theta)$.
  2. (ii)
    KL curvature expansion. Write $\delta=\theta-L_{n}$. The Taylor expansion of the Bernoulli KL around its minimiser $\theta=L_{n}$ gives
    $D_{\mathrm{KL}}(L_{n}\|\theta)=\frac{\delta^{2}}{2L_{n}(1-L_{n})}+O(\delta^{3}).$ (6)
    The quadratic coefficient is the Fisher information $\mathcal{I}(L_{n})=1/\bigl(L_{n}(1-L_{n})\bigr)$ evaluated at the empirical mean.
  3. (iii)
    First-moment coherence as linearisation. The martingale condition $\mathbb{E}[\theta\mid\mathcal{F}_{n}]=L_{n}$ constrains only the mean of the posterior (5): it pins the centre of the KL bowl but specifies nothing about the posterior’s spread or shape. First-moment martingale coherence is the linearisation of Bayesian updating: it retains the location of the posterior minimum but discards the curvature encoded in $\nabla^{2}D_{\mathrm{KL}}$ and all higher derivatives.
  4. (iv)
    Multi-step predictives require curvature. By Corollary 4.2, $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ depends on posterior moments up to order $k$. For $k=2$, the discrepancy between Bayes and plug-in equals the posterior variance $\sigma_{n}^{2}$ (equation (12)), which is governed by the quadratic term in (6). For $k\geq 3$, additional cumulants enter, corresponding to higher derivatives of the Sanov rate.
  5. (v)
    Derivative correspondence. Higher-order predictives are sensitive to higher-order features of the KL profile; in the Bernoulli case this is made explicit via the polynomial expansion $(1-\theta)^{k}=\sum_{j}\binom{k}{j}(-\theta)^{j}$ (Theorem 4.1). The moment hierarchy follows from this expansion, not from special structure of the Bernoulli model.
Proof.
(i) Since $\theta^{S_{n}}(1-\theta)^{n-S_{n}}=\exp\{S_{n}\log\theta+(n-S_{n})\log(1-\theta)\}$, adding and subtracting $S_{n}\log L_{n}+(n-S_{n})\log(1-L_{n})$ gives the exponent $n[L_{n}\log(L_{n}/\theta)+(1-L_{n})\log((1-L_{n})/(1-\theta))]=-n\,D_{\mathrm{KL}}(L_{n}\|\theta)$, plus the $\mathcal{F}_{n}$-measurable term $nH(L_{n})$.
(ii) Differentiating $f(\theta)=D_{\mathrm{KL}}(L_{n}\|\theta)=L_{n}\log(L_{n}/\theta)+(1-L_{n})\log((1-L_{n})/(1-\theta))$: $f^{\prime}(\theta)=-L_{n}/\theta+(1-L_{n})/(1-\theta)$, which vanishes at $\theta=L_{n}$. The second derivative is $f^{\prime\prime}(\theta)=L_{n}/\theta^{2}+(1-L_{n})/(1-\theta)^{2}$; at $\theta=L_{n}$ this equals $\mathcal{I}(L_{n})$. Taylor’s theorem gives (6).
(iii)–(v) follow from (i)–(ii) together with Theorem 4.1 and Corollary 4.2. ∎
3 Setup and notation
3.1 Exchangeable Bernoulli sequences
Let $(X_{i})_{i\geq 1}$ take values in $\{0,1\}$. Write $X_{1:n}=(X_{1},\ldots,X_{n})$ and $S_{n}=\sum_{i=1}^{n}X_{i}$.
Definition 3.1.
The sequence $(X_{i})_{i\geq 1}$ is exchangeable if for every $n\geq 1$ and every permutation $\sigma$ of $\{1,\ldots,n\}$, $(X_{1},\ldots,X_{n})\stackrel{{\scriptstyle d}}{{=}}(X_{\sigma(1)},\ldots,X_{\sigma(n)})$.
Theorem 3.2 (de Finetti).
Let $(X_{i})_{i\geq 1}$ be an exchangeable $\{0,1\}$-valued sequence. There exists a unique probability measure $\Pi$ on $[0,1]$ such that for every $n\geq 1$,
$\mathbb{P}(X_{1}=x_{1},\ldots,X_{n}=x_{n})=\int_{0}^{1}\theta^{s_{n}}(1-\theta)^{n-s_{n}}\,\mathrm{d}\Pi(\theta),\qquad s_{n}=\textstyle\sum_{i=1}^{n}x_{i}.$
Conditionally on $\theta$, the $X_{i}$ are i.i.d. $\mathrm{Bern}(\theta)$.
Proof.
Standard; see de Finetti (1937) and Hewitt–Savage (1955). The mixing measure is the law of $\lim_{n\to\infty}S_{n}/n$, identified via the ergodic decomposition of the exchangeable $\sigma$-algebra. ∎
The pair $(n,S_{n})$ is sufficient for the likelihood $\theta^{S_{n}}(1-\theta)^{n-S_{n}}$, so the posterior is
$\Pi(A\mid\mathcal{F}_{n})=\frac{\int_{A}\theta^{S_{n}}(1-\theta)^{n-S_{n}}\,\mathrm{d}\Pi(\theta)}{\int_{0}^{1}\theta^{S_{n}}(1-\theta)^{n-S_{n}}\,\mathrm{d}\Pi(\theta)},\qquad A\subseteq[0,1]\text{ measurable}.$ (7)
Remark 3.3 (Sufficiency).
The pair $(n,S_{n})$ is sufficient for $\theta$ in the Bernoulli likelihood. Every posterior quantity (moments, predictive probabilities, scoring-rule risks) reduces to an integral $\int f(\theta)\,\mathrm{d}\Pi(\theta\mid\mathcal{F}_{n})$ for appropriate $f$. The posterior (7) depends on the data only through $S_{n}$.
3.2 Martingale posteriors and the terminal value
Assumption 3.4.
$(\theta_{n})_{n\geq 0}$ is a sequence of $[0,1]$-valued, $\mathcal{F}_{n}$-measurable random variables satisfying the martingale condition (1).
Proposition 3.5 (Terminal value).
Under Assumption 3.4, $\theta_{n}\to\theta_{\infty}$ a.s. and in $L^{1}$, and $\mathbb{E}[\theta_{\infty}\mid\mathcal{F}_{n}]=\theta_{n}$ a.s. for all $n$. The random variable $\theta_{\infty}$ is the directing parameter.
Proof.
Since $(\theta_{n})$ is bounded in $[0,1]$, Doob’s forward convergence theorem gives a.s. convergence. Uniform boundedness gives $L^{1}$ convergence. The conditional expectation identity follows by passing to the limit in $\mathbb{E}[\theta_{m}\mid\mathcal{F}_{n}]=\theta_{n}$ ($m>n$); see Williams (1991), Theorem 11.5. ∎
Remark 3.6.
In the Bayesian framework with $\theta\sim\Pi$ and $X_{i}\mid\theta\sim\mathrm{Bern}(\theta)$, setting $\theta_{n}=\mathbb{E}[\theta\mid\mathcal{F}_{n}]$ yields a martingale with $\theta_{\infty}=\theta$ a.s. (Doob’s consistency theorem). The martingale condition is necessary for Bayesian coherence; the question is whether it is sufficient. Diaconis and Ylvisaker (1979) showed that conjugate priors are characterized by linearity of the posterior mean. Condition (1) is the sequential analogue.
Definition 3.7 (Predictive coherence).
A predictive system is $k$-step predictively coherent if, for every cylinder event of length $k$, the probability $\mathbb{P}(X_{n+1:n+k}=x_{1:k}\mid\mathcal{F}_{n})$ is correctly specified. It is fully predictively coherent if this holds for all finite $k$.
3.3 Martingale versus exchangeability
Proposition 3.8 (Martingale does not imply exchangeability).
There exist $[0,1]$-valued bounded martingales $(\theta_{n},\mathcal{F}_{n})$ that do not arise as posterior means $\mathbb{E}[\theta\mid\mathcal{F}_{n}]$ under any exchangeable Bernoulli law.
Proof.
Under exchangeability, the posterior mean (7) is a function of $(n,S_{n})$ alone. Let $\theta_{0}=1/2$ and define $\theta_{n}=\theta_{n-1}+2^{-n}(2X_{n}-1)$. By construction $(\theta_{n})$ is a bounded martingale, but $\theta_{n}$ depends on the order of observations, not merely $S_{n}$. Hence it cannot be a posterior mean under exchangeability. ∎
Definition 3.9 (Conditionally identically distributed).
A sequence $(X_{i})_{i\geq 1}$ is c.i.d. if for every $n\geq 1$ and every measurable $A$, $\mathbb{P}(X_{n+1}\in A\mid\mathcal{F}_{n})=\mathbb{P}(X_{n+2}\in A\mid\mathcal{F}_{n})$ a.s.
Proposition 3.10.
Every exchangeable Bernoulli sequence is c.i.d. The converse is false.
Proof.
Exchangeability $\Rightarrow$ c.i.d.: the conditional distribution of $X_{n+j}$ given $\mathcal{F}_{n}$ depends only on $(n,S_{n})$ and equals $\mathrm{Bern}(\mathbb{E}[\theta\mid\mathcal{F}_{n}])$ for all $j\geq 1$. For the converse, the martingale in Proposition 3.8 generates a c.i.d. sequence (future observations are conditionally $\mathrm{Bern}(\theta_{n})$) that is not exchangeable; see Berti, Pratelli, and Rigo (2004). ∎
For rates of convergence of predictive distributions under c.i.d., see Berti, Crimaldi, Pratelli and Rigo (2009).
Remark 3.11 (Role of c.i.d. in predictive coherence).
The c.i.d. property guarantees that the one-step conditional distribution $\mathbb{P}(X_{n+1}\in\cdot\mid\mathcal{F}_{n})$ is the same for all future indices $j\geq 1$. This is one-step predictive coherence (Definition 3.7 with $k=1$): knowing $\theta_{n}$ determines $\mathbb{P}(X_{n+j}=1\mid\mathcal{F}_{n})=\theta_{n}$ for every $j$. However, c.i.d. does not imply that future observations are conditionally independent given $\mathcal{F}_{n}$; that requires knowledge of the full conditional law of $\theta_{\infty}$. For $k\geq 2$, the joint predictive $\mathbb{P}(X_{n+1}=x_{1},X_{n+2}=x_{2}\mid\mathcal{F}_{n})$ depends on the posterior variance through (12), which is not determined by the c.i.d. property alone. Thus c.i.d. $\Rightarrow$ 11-step coherence, but c.i.d. $\not\Rightarrow$ $k$-step coherence for $k\geq 2$, paralleling Theorem 6.3 at the process level.
4 The predictive moment hierarchy
4.1 Moment representation
Theorem 4.1 (Moment representation).
Let $(X_{i})_{i\geq 1}$ be exchangeable Bernoulli with mixing measure $\Pi$. For any $n\geq 0$, $k\geq 1$, and pattern $x_{1:k}\in\{0,1\}^{k}$ with $s=\sum_{i=1}^{k}x_{i}$,
$\mathbb{P}(X_{n+1:n+k}=x_{1:k}\mid\mathcal{F}_{n})=\mathbb{E}\bigl[\theta^{s}(1-\theta)^{k-s}\mid\mathcal{F}_{n}\bigr].$ (8)
In particular, the $k$-step run probability is
$\mathbb{P}(X_{n+1}=\cdots=X_{n+k}=0\mid\mathcal{F}_{n})=\mathbb{E}\bigl[(1-\theta)^{k}\mid\mathcal{F}_{n}\bigr].$ (9)
Proof.
By Theorem 3.2, $X_{n+1},\ldots,X_{n+k}$ are conditionally i.i.d. $\mathrm{Bern}(\theta)$ and conditionally independent of $\mathcal{F}_{n}$. Hence $\mathbb{P}(X_{n+1:n+k}=x_{1:k}\mid\mathcal{F}_{n},\theta)=\theta^{s}(1-\theta)^{k-s}$. Integrate over $\Pi(\,\mathrm{d}\theta\mid\mathcal{F}_{n})$ by the tower property. ∎
Corollary 4.2 (Moment expansion).
For any $k\geq 1$,
$\mathbb{E}\bigl[(1-\theta)^{k}\mid\mathcal{F}_{n}\bigr]=\sum_{j=0}^{k}\binom{k}{j}(-1)^{j}\,\mathbb{E}[\theta^{j}\mid\mathcal{F}_{n}].$ (10)
The $k$-step run probability depends on posterior moments up to order $k$.
Proof.
Apply the binomial theorem to $(1-\theta)^{k}=\sum_{j=0}^{k}\binom{k}{j}(-1)^{j}\theta^{j}$ and take conditional expectations. ∎
Each additional prediction step introduces exactly one new posterior moment. The moment hierarchy thereby induces a strictly increasing filtration of predictive $\sigma$-algebras indexed by horizon length.
4.2 Injectivity
Theorem 4.3 (Injectivity).
The mapping $\Pi(\cdot\mid\mathcal{F}_{n})\longmapsto\bigl\{\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]:k\geq 1\bigr\}$ is injective. The full sequence of $k$-step run probabilities uniquely determines the posterior $\Pi(\cdot\mid\mathcal{F}_{n})$.
Proof.
By Corollary 4.2, knowing $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ for all $k\geq 1$ determines the moment sequence $(\mathbb{E}[\theta^{j}\mid\mathcal{F}_{n}])_{j\geq 1}$ via Möbius inversion (Proposition 4.6). On $[0,1]$, the moment sequence uniquely determines the measure (Hausdorff moment theorem; Theorem A.1 in the Appendix). ∎
Remark 4.4 (Polynomial density).
Theorem 4.3 also follows from Weierstrass approximation: polynomials are dense in $C[0,1]$, so the moment sequence determines all expectations $\mathbb{E}[f(\theta)\mid\mathcal{F}_{n}]$ for $f\in C[0,1]$, hence the measure. Injectivity relies on the compact support $[0,1]$; on $\mathbb{R}$, the Stieltjes moment problem may have non-unique solutions (e.g., the log-normal distribution).
Remark 4.5 (Comparison with the real line).
On $[0,1]$, the moment problem is determinate: a probability measure is uniquely determined by its moment sequence (Theorem A.1). This is the mechanism behind Theorem 4.3. On $\mathbb{R}$, the Hamburger moment problem is indeterminate in general: distinct measures can share all moments. A classical example is the log-normal: $X\sim\mathrm{LogNormal}(0,1)$ has the same moments as the family $X\cdot(1+a\sin(2\pi\log X))$ for $|a|\leq 1$. The moment hierarchy established in Section 4 therefore provides a complete characterization of the posterior only because $\theta$ takes values in the compact interval $[0,1]$. For parameters in $\mathbb{R}^{d}$, analogous results would require additional conditions such as exponential moment bounds (Carleman’s condition) or restriction to exponential families with compact sufficient statistics.
4.3 Binomial inversion
Proposition 4.6 (Möbius inversion).
Let $m_{k}=\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ and $\mu_{j}=\mathbb{E}[\theta^{j}\mid\mathcal{F}_{n}]$ with $\mu_{0}=1$. Then
$\mu_{j}=\sum_{\ell=0}^{j}\binom{j}{\ell}(-1)^{\ell}\,m_{\ell},\qquad j\geq 0.$ (11)
The matrix relating $(\mu_{0},\ldots,\mu_{k})$ to $(m_{0},\ldots,m_{k})$ is upper-triangular with diagonal entries $\pm 1$, hence invertible.
Proof.
Expand $\theta^{j}=(1-(1-\theta))^{j}=\sum_{\ell=0}^{j}\binom{j}{\ell}(-1)^{\ell}(1-\theta)^{\ell}$ and take conditional expectations. ∎
Example 4.7 (Inversion at $j=2$).
From the forward relation: $m_{2}=1-2\mu_{1}+\mu_{2}$. By Möbius inversion (11): $\mu_{2}=m_{0}-2m_{1}+m_{2}$, i.e., $\mu_{2}=1-2(1-\mu_{1})+m_{2}=2\mu_{1}-1+m_{2}$. Substituting $m_{2}=(1-\mu_{1})^{2}+\sigma_{n}^{2}$ gives $\mu_{2}=2\mu_{1}-1+(1-\mu_{1})^{2}+\sigma_{n}^{2}=\mu_{1}^{2}+\sigma_{n}^{2}$, recovering $\sigma_{n}^{2}=\mu_{2}-\mu_{1}^{2}$ as expected. This illustrates the triangular structure: the $j$-th moment and the $j$-step run probability determine each other given lower-order quantities.
4.4 Explicit expansion for small $k$
Write $\mu_{j}=\mathbb{E}[\theta^{j}\mid\mathcal{F}_{n}]$.
For $k=1$: $\mathbb{E}[(1-\theta)\mid\mathcal{F}_{n}]=1-\mu_{1}$.
For $k=2$:
$\mathbb{E}[(1-\theta)^{2}\mid\mathcal{F}_{n}]=(1-\mu_{1})^{2}+\mathrm{Var}(\theta\mid\mathcal{F}_{n}).$ (12)
The two-step predictive depends on the posterior variance, which is not determined by the posterior mean. The discrepancy between Bayes and plug-in equals $\mathrm{Var}(\theta\mid\mathcal{F}_{n})$ exactly.
For $k=3$: $\mathbb{E}[(1-\theta)^{3}\mid\mathcal{F}_{n}]=1-3\mu_{1}+3\mu_{2}-\mu_{3}$.
For $k=4$:
$\mathbb{E}[(1-\theta)^{4}\mid\mathcal{F}_{n}]$ $=1-4\mu_{1}+6\mu_{2}-4\mu_{3}+\mu_{4}$
$=(1-\mu_{1})^{4}+6(1-\mu_{1})^{2}\sigma_{n}^{2}+\bigl(4\kappa_{3}(1-\mu_{1})+\kappa_{4}+3\sigma_{n}^{4}\bigr),$ (13)
where $\kappa_{3}=\mathbb{E}[(\theta-\mu_{1})^{3}\mid\mathcal{F}_{n}]$ is the third central moment and $\kappa_{4}=\mathbb{E}[(\theta-\mu_{1})^{4}\mid\mathcal{F}_{n}]-3\sigma_{n}^{4}$ is the fourth cumulant (excess kurtosis contribution).
Each additional step introduces one new moment: $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ cannot be expressed as a function of $\mu_{1},\ldots,\mu_{k-1}$ alone, since the coefficient of $(-1)^{k}\mu_{k}$ in (10) is $\binom{k}{k}=1\neq 0$.
5 KL geometry and the moment hierarchy
This section develops the KL geometry underlying the moment hierarchy, adding two structural results (Theorem 5.1 and Proposition 5.2) and a diagram.
For Bernoulli, $D_{\mathrm{KL}}(p\|\cdot)$ has Fisher information $\mathcal{I}(\theta)=1/(\theta(1-\theta))$. Since $m_{n}$ minimises $D_{\mathrm{KL}}(m_{n}\|\cdot)$, its gradient vanishes at $m_{n}$; the curvature there is $\mathcal{I}(m_{n})=1/(m_{n}(1-m_{n}))$. The posterior variance satisfies $\sigma_{n}^{2}\approx m_{n}(1-m_{n})/n$ (Bernstein–von Mises), the reciprocal of this curvature. Predictively, the hierarchy mirrors a derivative hierarchy:
$\text{$k=1$: }\nabla D_{\mathrm{KL}}\big|_{m_{n}}=0,\qquad\text{$k\geq 2$: }\nabla^{k}D_{\mathrm{KL}}\big|_{m_{n}}\neq 0.$ (14)
The one-step predictive $1-m_{n}$ depends only on the zeroth-order term; the two-step predictive is the first to require curvature (equation (12)).
5.1 The Sanov Gap
Theorem 5.1 (Sanov Gap).
Let $m\in(0,1)$.
  1. (i)
    First-order indistinguishability. For any two posteriors $\nu_{1},\nu_{2}$ on $[0,1]$ with $\int\theta\,\mathrm{d}\nu_{i}=m$,
    $\frac{\partial}{\partial\theta}D_{\mathrm{KL}}(m\|\theta)\Big|_{\theta=m}=0,$
    so they are indistinguishable at first order in the Sanov expansion: the linear approximation to the posterior log-likelihood is the same for both.
  2. (ii)
    Second-order separation. The quadratic correction to the $k$-step predictive equals the posterior variance:
    $\mathbb{E}\bigl[(1-\theta)^{k}\mid\mathcal{F}_{n}\bigr]-(1-m)^{k}=\tfrac{k(k-1)}{2}(1-\xi)^{k-2}\sigma^{2},\quad\sigma^{2}=\int(\theta-m)^{2}\,\mathrm{d}\nu.$
    Distinct values of $\sigma^{2}$ yield distinct $k$-step predictives for all $k\geq 2$.
  3. (iii)
    Non-pathological gap. For any $m\in(0,1)$ and $k\geq 2$, the $k$-step predictive ranges over a nondegenerate interval as $\nu$ ranges over posteriors with mean $m$; the gap is not a boundary phenomenon (Remark 6.6).
Proof.
(i) $D_{\mathrm{KL}}(m\|\theta)$ is minimised at $\theta=m$, so its derivative vanishes there (Theorem 2.1(ii)). (ii) is equation (15) (Proposition 6.4). (iii) follows from Lemma 6.1 and Remark $6.$ ∎
Theorem 5.1 is the geometric source of the non-uniqueness in Theorem 6.3: posteriors sharing the same mean are first-order equivalent in the Sanov expansion but separate at second order via $\sigma^{2}$.
5.2 Moderate deviations and the second moment as first nontrivial Sanov term
Proposition 5.2 (Second moment as leading Sanov curvature).
Under the moderate-deviations scaling ($a_{n}/\sqrt{n}\to\infty$, $a_{n}=o(\sqrt{n})$), the posterior satisfies a moderate-deviation principle with quadratic rate $\frac{1}{2}(\theta-\theta_{0})^{2}\mathcal{I}(\theta_{0})$ (Eichelsbacher and Ganesh, 2002). Consequently:
  1. (i)
    The posterior variance satisfies $\sigma_{n}^{2}=m_{n}(1-m_{n})/n+O(n^{-2})$, recovering the Bernstein–von Mises rate.
  2. (ii)
    The second moment is the first nontrivial curvature term: it corresponds to $\nabla^{2}D_{\mathrm{KL}}|_{m_{n}}$ and cannot be recovered from the zeroth- and first-order terms.
  3. (iii)
    Specifying only the posterior mean (first-moment coherence) is equivalent to retaining the zeroth-order Sanov term; the second moment requires the quadratic term; the $k$-th moment requires the $k$-th derivative.
Figure 1 plots the Sanov rate $D_{\mathrm{KL}}(m_{n}\|\theta)$ for $m_{n}=0.4$, its quadratic approximation, and the resulting posterior density.
[…]
Figure 1: Sanov geometry ($m_{n}=0.4$, $n=12$). Blue: the Sanov rate $D_{\mathrm{KL}}(m_{n}\|\theta)$, which governs posterior shape through (5). Dashed red: its quadratic (Fisher information) approximation, retaining only the second-order term. Green: the resulting posterior density. The dotted vertical at $m_{n}$ marks what first-moment information identifies. One-step prediction requires only this location; two-step additionally requires the curvature $\sigma_{n}^{2}$ (equation (12)); $k$-step prediction for $k\geq 3$ requires further moments.
6 Insufficiency of first-moment coherence
6.1 Non-identification and moment insufficiency
Lemma 6.1 (Non-identification).
For any $m\in(0,1)$ and any integer $k\geq 2$, there exist probability measures $\nu_{1},\nu_{2}$ on $[0,1]$ with $\int\theta\,\mathrm{d}\nu_{1}=\int\theta\,\mathrm{d}\nu_{2}=m$ but $\int(1-\theta)^{k}\,\mathrm{d}\nu_{1}\neq\int(1-\theta)^{k}\,\mathrm{d}\nu_{2}$.
Proof.
Let $\nu_{1}=\delta_{m}$, so $\int(1-\theta)^{k}\,\mathrm{d}\nu_{1}=(1-m)^{k}$. Choose $0<a<m<b<1$ and set $\nu_{2}=\lambda\delta_{a}+(1-\lambda)\delta_{b}$ with $\lambda=(b-m)/(b-a)\in(0,1)$, ensuring $\int\theta\,\mathrm{d}\nu_{2}=m$. Since $\theta\mapsto(1-\theta)^{k}$ is strictly convex for $k\geq 2$, Jensen’s inequality gives $\int(1-\theta)^{k}\,\mathrm{d}\nu_{2}>(1-m)^{k}=\int(1-\theta)^{k}\,\mathrm{d}\nu_{1}$. ∎
Remark 6.2 (Extension to general functionals).
Lemma 6.1 extends to any non-affine continuous $f:[0,1]\to\mathbb{R}$: the mean does not determine $\mathbb{E}[f(\theta)\mid\mathcal{F}_{n}]$ (Appendix B).
6.2 Numerical illustration
At $n=0$, take $\Pi_{1}=\mathrm{Beta}(2,2)$ and $\Pi_{2}=\mathrm{Beta}(1,1)$. Both have mean $1/2$. Since $\mathbb{E}_{\mathrm{Beta}(a,b)}[\theta^{2}]=a(a+1)/\bigl((a+b)(a+b+1)\bigr)$:
$\mathbb{E}_{\Pi_{1}}[(1-\theta)^{2}]$ $=\tfrac{3}{10},\qquad\mathbb{E}_{\Pi_{2}}[(1-\theta)^{2}]=\tfrac{1}{3}.$
Equal prior means; different two-step predictives. The variances are $1/20$ and $1/12$ respectively. These priors do not maintain equal posterior means for $n\geq 1$: the posterior mean $(a+S_{n})/(a+b+n)$ depends on $a+b$.
Theorem 6.3 (Moment insufficiency).
Let $k\geq 2$, $n\geq 0$, and $m_{n}\coloneqq\mathbb{E}[\theta\mid\mathcal{F}_{n}]\in(0,1)$. The mapping $m_{n}\mapsto\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ is set-valued: distinct posteriors sharing mean $m_{n}$ yield distinct $k$-step predictives.
Proof.
By Lemma 6.1 and Theorem 4.1: any two measures $\nu_{1},\nu_{2}$ on $[0,1]$ with equal mean but different $k$-th moments (for $k\geq 2$) serve as de Finetti mixing measures with the same posterior mean $m_{n}$ but different $k$-step predictives. ∎
6.3 Quantitative bound
Proposition 6.4 (Second-order discrepancy).
Let $k\geq 2$ and let $\Pi(\cdot\mid\mathcal{F}_{n})$ have mean $m_{n}$ and variance $\sigma_{n}^{2}$. Then
$\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-m_{n})^{k}=\frac{k(k-1)}{2}(1-\xi_{n})^{k-2}\,\sigma_{n}^{2}$ (15)
for some $\xi_{n}\in[0,1]$. In particular,
$0<\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-m_{n})^{k}\leq\frac{k(k-1)}{2}\,\sigma_{n}^{2}.$ (16)
Proof.
Set $f(\theta)=(1-\theta)^{k}$. By Taylor’s theorem with Lagrange remainder around $m_{n}$: $f(\theta)=f(m_{n})+f^{\prime}(m_{n})(\theta-m_{n})+\tfrac{1}{2}f^{\prime\prime}(\xi)(\theta-m_{n})^{2}$. Take conditional expectations and use $\mathbb{E}[\theta-m_{n}\mid\mathcal{F}_{n}]=0$:
$\mathbb{E}[f(\theta)\mid\mathcal{F}_{n}]-f(m_{n})=\tfrac{1}{2}\mathbb{E}[f^{\prime\prime}(\xi)(\theta-m_{n})^{2}\mid\mathcal{F}_{n}].$
Since $f^{\prime\prime}(\theta)=k(k-1)(1-\theta)^{k-2}\geq 0$ on $[0,1]$ (strict convexity) and $\sup f^{\prime\prime}=k(k-1)$, the mean value theorem for integrals gives (15). The lower bound is strict when $\sigma_{n}^{2}>0$ by strict convexity. ∎
Corollary 6.5 (Plug-in as Dirac).
The plug-in predictive $\hat{p}_{k}^{\mathrm{plug}}\coloneqq(1-\theta_{n})^{k}$ equals $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ under $\delta_{\theta_{n}}$. For $k\geq 2$, it systematically underestimates the run probability by at most $\frac{k(k-1)}{2}\sigma_{n}^{2}$. At $k=2$, the discrepancy equals $\sigma_{n}^{2}$ exactly by (12).
Remark 6.6 (Strength of non-identification).
The non-identification in Lemma 6.1 is not a pathological boundary phenomenon. For any $m\in(0,1)$ and any $k\geq 2$, the set of posteriors with mean $m$ is an infinite-dimensional convex set, and the $k$-step predictive $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ takes all values in the interval $\bigl[(1-m)^{k},\,(1-m)^{k}+\tfrac{k(k-1)}{2}m(1-m)\bigr)$ as the posterior ranges over distributions with mean $m$. The lower bound is attained by $\delta_{m}$ (plug-in) and the upper bound is approached by two-point distributions concentrating near 0 and $1$. At $k=2$ with $m=1/2$, the Bayes predictive ranges from $1/4$ (degenerate) to $1/2$ (symmetric Bernoulli on $\{0,1\}$), a factor of two.
Remark 6.7 (Scope).
Theorem 6.3 does not assert that the martingale posterior framework is incompatible with Bayesian inference. Every Bayesian posterior mean is a martingale (Remark 3.6). The theorem asserts that condition (1) alone does not determine the full posterior. A specific likelihood, a named prior, or explicit higher-moment constraints are needed for multi-step predictive coherence (Definition 3.7).
7 Decision-theoretic consequence
Let $\delta$ be a predictive rule producing $\hat{P}_{\delta}(\cdot\mid\mathcal{F}_{n})$ on $\{0,1\}^{k}$.
Definition 7.1.
A scoring rule $S(\hat{P},x)$ assigns a loss to predictive $\hat{P}$ when outcome $x$ is realized. It is strictly proper if $\mathbb{E}_{P}[S(\hat{P},X)]$ is uniquely minimized at $\hat{P}=P$. The risk is $R(\delta,\theta)=\mathbb{E}_{\theta}[S(\hat{P}_{\delta},X_{n+1:n+k})]$. A rule is admissible if no rule uniformly dominates it in risk.
The log score $S(\hat{P},x)=-\log\hat{P}(x)$ and Brier score $S(\hat{P},x)=\sum_{y}(\mathbf{1}[y=x]-\hat{P}(y))^{2}$ are both strictly proper (Gneiting and Raftery, 2007).
7.1 Complete class theorem
The following result is classical and included for completeness; see Blackwell and Girshick (1954) and Wald (1950).
Theorem 7.2 (Blackwell–Girshick).
Let $S$ be a strictly proper scoring rule with parameter space $[0,1]$ and risk continuous in $\theta$. Then:
  1. (a)
    Every Bayes rule is admissible.
  2. (b)
    Every admissible rule is Bayes or a limit of Bayes rules.
The unique Bayes-optimal predictive under prior $\Pi$ is the posterior predictive $\hat{P}^{\mathrm{Bayes}}(x\mid\mathcal{F}_{n})=\mathbb{E}[\theta^{s}(1-\theta)^{k-s}\mid\mathcal{F}_{n}]$.
Proof.
Part (a) is standard. Part (b) follows from Blackwell–Girshick (1954) (compact parameter space, continuous risk). Uniqueness follows from strict propriety (Gneiting and Raftery, 2007). ∎
7.2 Plug-in dominance
Proposition 7.3 (Plug-in is strictly dominated).
Let $\Pi$ have full support on $[0,1]$ and let $S$ be any strictly proper scoring rule. For $k\geq 2$, the plug-in rule $\delta^{\mathrm{plug}}$ is strictly dominated by $\delta^{\mathrm{Bayes}}$: for every $\theta\in(0,1)$, $R(\delta^{\mathrm{Bayes}},\theta)<R(\delta^{\mathrm{plug}},\theta)$.
Proof.
We show that plug-in and Bayes predictives differ almost surely, then apply strict propriety.
Fix $X_{1:n}$ with $\mathrm{Var}(\theta\mid\mathcal{F}_{n})>0$ (which holds $\mathbb{P}_{\theta}$-a.s. since $\Pi$ has full support). The Bayes predictive is $\hat{P}^{\mathrm{Bayes}}(x)=\mathbb{E}[\theta^{s}(1-\theta)^{k-s}\mid\mathcal{F}_{n}]$; the plug-in is $\hat{P}^{\mathrm{plug}}(x)=\theta_{n}^{s}(1-\theta_{n})^{k-s}$. By Corollary 6.5, for $k\geq 2$ these differ on a set of positive probability whenever the posterior is non-degenerate. Since the Bayes predictive equals the conditional distribution of the future block given $\mathcal{F}_{n}$, strict propriety implies that $\hat{P}^{\mathrm{Bayes}}$ strictly minimises expected score. Hence the plug-in’s expected score is strictly higher. Integrating over $\mathcal{F}_{n}$ under $\mathbb{P}_{\theta}$ preserves the strict inequality. ∎
The dominance is strict whenever $\mathrm{Var}(\theta\mid\mathcal{F}_{n})>0$. For a geometric interpretation of admissibility via risk-set boundaries and supporting hyperplanes, and its extension to sequential, coverage, and approachability criteria, see Polson and Zantedeschi (2026).
Corollary 7.4 (Inadmissibility).
The plug-in rule is inadmissible for $k$-step prediction with $k\geq 2$ under any strictly proper scoring rule.
Proof.
Immediate from Proposition 7.3: the Bayes rule strictly dominates the plug-in whenever $\Pi$ has full support. ∎
Corollary 7.5 (Mean-only rules).
Let $k\geq 2$. Any predictive rule depending on $\mathcal{F}_{n}$ only through $\theta_{n}$ is either a Bayes rule under a degenerate prior, or outside the Bayes class and hence dominable.
Proof.
By Theorem 6.3, two posteriors sharing the same mean generically yield different Bayes-optimal predictives for $k\geq 2$. A mean-only rule assigns the same predictive to both; it cannot be Bayes-optimal for both simultaneously. ∎
Remark 7.6 (Log-score risk gap).
At $k=2$, the Bayes predictive assigns run probability $p_{B}=\mathbb{E}[(1-\theta)^{2}\mid\mathcal{F}_{n}]=(1-m_{n})^{2}+\sigma_{n}^{2}$ and the plug-in assigns $p_{P}=(1-m_{n})^{2}$. The KL divergence between these Bernoulli predictives satisfies $D_{\mathrm{KL}}(\text{Bayes}\|\text{plug-in})\geq\sigma_{n}^{4}/(2\ln 2)$ by Pinsker’s inequality, giving a lower bound on the risk gap that is quadratic in the posterior variance.
Example 7.7 (Numerical risk gap).
Continue the setting of Table 3: $n=5$, $S_{5}=2$, Jeffreys prior, so $m_{5}=5/12$ and $\sigma_{5}^{2}\approx 0.035$.
At $k=2$, the Bayes predictive for the all-zeros run is $p_{B}=(b_{5})_{2}/(N_{5})_{2}=0.375$ and the plug-in is $p_{P}=(1-m_{5})^{2}\approx 0.340$. The log-score risk gap for predicting the binary event $Y=\mathbf{1}[X_{6}=X_{7}=0]$ is
$\Delta R_{\log}$ $=\mathbb{E}_{Y}[-\log p_{P}(Y)]-\mathbb{E}_{Y}[-\log p_{B}(Y)]$
$=p_{B}\log\frac{p_{B}}{p_{P}}+(1-p_{B})\log\frac{1-p_{B}}{1-p_{P}}$
$\approx 0.375\log\frac{0.375}{0.340}+0.625\log\frac{0.625}{0.660}\approx 0.0026.$
This KL divergence, while small in absolute terms, is strictly positive and cannot be eliminated by any mean-only adjustment. At $k=4$ the divergence grows to approximately $0.021$, an order of magnitude larger. Accumulated over a sequential prediction task with $T$ rounds, the total excess log-loss is $\Omega(T\sigma_{n}^{4})$, non-negligible when $n$ is moderate relative to $T$.
Remark 7.8 (Brier score risk gap).
Under the Brier score $S(\hat{p},y)=(\hat{p}-y)^{2}$ for predicting the binary event $Y=\mathbf{1}[X_{n+1}=\cdots=X_{n+k}=0]$, the risk gap between plug-in and Bayes is $(p_{B}-p_{P})^{2}$ to leading order. At $k=2$, this equals $\sigma_{n}^{4}$. The qualitative conclusion (plug-in is strictly dominated whenever $\sigma_{n}^{2}>0$) holds under all strictly proper scoring rules.
Under the log score, the loss from using $q$ instead of $p$ equals $D_{\mathrm{KL}}(p\|q)$ (Cover and Thomas, 2006): KL is simultaneously predictive regret and the canonical divergence between distributions. The Bayes predictive minimizes expected loss, and the gap to any plug-in baseline is a KL term (Proposition 7.3, Remark 7.6).
8 Hill’s $A_{(n)}$ as a positive example
Hill’s $A_{(n)}$ (Hill, 1968, 1988), in the Bernoulli case, is the posterior predictive under the Jeffreys prior $\mathrm{Beta}(\tfrac{1}{2},\tfrac{1}{2})$.
Proposition 8.1 (Hill/Jeffreys one-step predictive).
Let $\theta\sim\mathrm{Beta}(\tfrac{1}{2},\tfrac{1}{2})$ and $X_{i}\mid\theta\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Bern}(\theta)$. The posterior is $\mathrm{Beta}(\tfrac{1}{2}+S_{n},\,\tfrac{1}{2}+n-S_{n})$, and $\mathbb{P}(X_{n+1}=1\mid\mathcal{F}_{n})=(S_{n}+\tfrac{1}{2})/(n+1)$.
Proof.
The posterior mean of $\mathrm{Beta}(a,b)$ is $a/(a+b)$. With $a=\tfrac{1}{2}+S_{n}$ and $b=\tfrac{1}{2}+n-S_{n}$, the result follows by Theorem 4.1 with $k=1$. ∎
Proposition 8.2 ($k$-step run probabilities).
Under the same model, with $a_{n}=S_{n}+\tfrac{1}{2}$, $b_{n}=n-S_{n}+\tfrac{1}{2}$, $N_{n}=a_{n}+b_{n}=n+1$, and $(x)_{k}=x(x+1)\cdots(x+k-1)$ the rising factorial,
$\mathbb{P}(X_{n+1}=\cdots=X_{n+k}=0\mid\mathcal{F}_{n})=\prod_{j=0}^{k-1}\frac{b_{n}+j}{N_{n}+j}.$
Proof.
By Theorem 4.1 and the Beta moment formula: $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]=B(a_{n},b_{n}+k)/B(a_{n},b_{n})=(b_{n})_{k}/(N_{n})_{k}$. ∎
8.1 Full moment recovery
Under $\mathrm{Beta}(a_{n},b_{n})$, all posterior moments are determined:
$\mathbb{E}[\theta^{j}\mid\mathcal{F}_{n}]=\frac{(a_{n})_{j}}{(N_{n})_{j}}=\prod_{i=0}^{j-1}\frac{a_{n}+i}{N_{n}+i},\qquad j\geq 1.$ (17)
Hill’s rule specifies every conditional moment as an explicit function of $(n,S_{n})$. It is a Bayes rule and therefore admissible (Theorem 7.2(a)).
The first three moments with $N_{n}=n+1$ are:
$\mu_{1}$ $=\frac{a_{n}}{N_{n}}=\frac{S_{n}+\tfrac{1}{2}}{n+1},$ (18)
$\mu_{2}$ $=\frac{a_{n}(a_{n}+1)}{N_{n}(N_{n}+1)}=\frac{(S_{n}+\tfrac{1}{2})(S_{n}+\tfrac{3}{2})}{(n+1)(n+2)},$ (19)
$\mu_{3}$ $=\frac{a_{n}(a_{n}+1)(a_{n}+2)}{N_{n}(N_{n}+1)(N_{n}+2)}=\frac{(S_{n}+\tfrac{1}{2})(S_{n}+\tfrac{3}{2})(S_{n}+\tfrac{5}{2})}{(n+1)(n+2)(n+3)}.$ (20)
Each is a rational function of $(n,S_{n})$; the posterior variance is $\sigma_{n}^{2}=\mu_{2}-\mu_{1}^{2}=a_{n}b_{n}/(N_{n}^{2}(N_{n}+1))$.
8.2 Plug-in versus Bayes comparison
Table 3 compares the plug-in and Bayes (Jeffreys) $k$-step run probabilities $\mathbb{P}(X_{n+1:n+k}=0\mid\mathcal{F}_{n})$ for $n=5$, $S_{5}=2$ ($m_{5}=5/12\approx 0.417$).
Table 3: Plug-in vs. Bayes (Jeffreys) $k$-step run probabilities $\mathbb{P}(X_{n+1:n+k}=0\mid\mathcal{F}_{n})$, $n=5$, $S_{5}=2$.
$k$ Plug-in $(1-m_{5})^{k}$ Bayes $(b_{5})_{k}/(N_{5})_{k}$ Relative gap
2 0.340 0.375 9.3%
3 0.199 0.258 23.0%
4 0.116 0.186 37.8%
The Jeffreys posterior variance is $\sigma_{5}^{2}=a_{5}b_{5}/(N_{5}^{2}(N_{5}+1))=(2.5)(3.5)/(252)\approx 0.035$, matching the $k=2$ discrepancy by (12).
To verify: with $a_{5}=2.5$, $b_{5}=3.5$, $N_{5}=6$, the Bayes predictives are $(b_{5})_{k}/(N_{5})_{k}$:
$k=2:$ $\quad(3.5)(4.5)/\bigl((6)(7)\bigr)=0.375,$
$k=3:$ $\quad(3.5)(4.5)(5.5)/\bigl((6)(7)(8)\bigr)\approx 0.258,$
$k=4:$ $\quad(3.5)(4.5)(5.5)(6.5)/\bigl((6)(7)(8)(9)\bigr)\approx 0.186.$
The plug-in values are $(7/12)^{k}$: $0.340$, $0.199$, $0.116$. The relative gap grows from $9.3\%$ ($k=2$) to $37.8\%$ ($k=4$), reflecting accumulating dependence on higher-order moments: for $k\geq 3$, the second-order Taylor bound $\frac{k(k-1)}{2}\sigma_{n}^{2}$ underestimates the true discrepancy.
Remark 8.3 (Lane–Sudderth caveat).
Hill’s $A_{(n)}$ was formulated nonparametrically. Lane and Sudderth (1984) showed that in nonparametric settings, conglomerability can fail. In the Bernoulli case, $\mathrm{Beta}(\tfrac{1}{2},\tfrac{1}{2})$ is a proper prior and no such issue arises.
Remark 8.4 (Red/black card prediction).
In the standard deck-prediction problem, one-step prediction depends only on the posterior mean (martingale coherence), but $k$-step run probabilities require $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$, hence all posterior moments up to order $k$ (Corollary 4.2). The curvature of $(1-\theta)^{k}$ makes the variance contribution non-negligible whenever the remaining composition deviates from $1/2$.
9 Asymptotic reconciliation
The moment insufficiency of Section 6 is a finite-sample phenomenon.
Proposition 9.1 (Asymptotic agreement).
Let $k\geq 1$ be fixed and let $\sigma_{n}^{2}=\mathrm{Var}(\theta\mid\mathcal{F}_{n})\to 0$ a.s. Then
$\bigl|\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-\mathbb{E}[\theta\mid\mathcal{F}_{n}])^{k}\bigr|\leq\tfrac{k(k-1)}{2}\,\sigma_{n}^{2}\to 0\quad\text{a.s.}$
Proof.
By Proposition 6.4. ∎
Under $\mathrm{Beta}(a,b)$ prior, the posterior variance is
$\sigma_{n}^{2}=\frac{(a+S_{n})(b+n-S_{n})}{(a+b+n)^{2}(a+b+n+1)}.$ (21)
Since $S_{n}/n\to\theta_{0}$ a.s., $\sigma_{n}^{2}\sim\theta_{0}(1-\theta_{0})/n$, confirming $\sigma_{n}^{2}=O(1/n)$. The following proposition gives a uniform-in-$k$ scaling result.
Proposition 9.2 (Uniform bound).
Let $K_{n}=o(\sqrt{n})$. Then $\sup_{1\leq k\leq K_{n}}\bigl|\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-m_{n})^{k}\bigr|\to 0$ a.s.
Proof.
By Proposition 6.4, $\sup_{k\leq K_{n}}\frac{k(k-1)}{2}\sigma_{n}^{2}\leq\frac{K_{n}^{2}}{2}\cdot O(1/n)\to 0$ when $K_{n}=o(\sqrt{n})$. ∎
Proposition 9.3 (Bernstein–von Mises rate).
Let $\Pi$ have a continuous, positive density near $\theta_{0}\in(0,1)$. For fixed $k$,
$\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]-(1-m_{n})^{k}=\frac{k(k-1)}{2}(1-\theta_{0})^{k-2}\cdot\frac{\theta_{0}(1-\theta_{0})}{n}+O(n^{-2})\quad\text{a.s.}$
The relative discrepancy is $O(1/n)$.
Proof.
By the Bernstein–von Mises theorem, $\sigma_{n}^{2}=\theta_{0}(1-\theta_{0})/n+O(n^{-2})$ a.s. Substituting into Proposition 6.4 and noting $\xi_{n}\to\theta_{0}$ a.s. gives the expansion. ∎
Remark 9.4 (Finite-sample relevance).
The discrepancy vanishes as $n\to\infty$ for fixed $k$. When the prediction horizon grows with $n$, Proposition 9.2 requires $k=o(\sqrt{n})$; for $k\geq C\sqrt{n}$, the bound $\frac{k(k-1)}{2}\sigma_{n}^{2}$ remains $O(1)$.
10 Structural hierarchy
$\mathbb{E}[\theta_{\infty}\mid\mathcal{F}_{n}]$$\bigl(\mathbb{E}[\theta_{\infty}^{j}\mid\mathcal{F}_{n}]\bigr)_{j\leq J}$$\mathcal{L}(\theta_{\infty}\mid\mathcal{F}_{n})$$k=1$ $k\leq J predictive$sall $k$-step predictives
Figure 2: The predictive moment hierarchy. Bottom: the martingale / mean-only tier (cf. linear Bayes) constrains only the first moment and determines one-step predictives. Middle: specifying $J$ moments (Goldstein (1975a) prevision tier) determines $k$-step predictives for $k\leq J$ but not for $k>J$ (Theorem 6.3). Top: the full conditional law determines all $k$-step predictives (predictive completeness; Theorem 10.3).
10.1 Hierarchy of predictive frameworks
Table 4 summarizes the relationship between the structure imposed by a predictive framework and the moments it determines.
Table 4: Structural hierarchy for exchangeable Bernoulli sequences.
Framework Structure imposed Moments determined
Martingale posterior $\mathbb{E}[\theta_{\infty}\mid\mathcal{F}_{n}]=\theta_{n}$ 1st moment
$j$-moment specification $\mu_{1},\ldots,\mu_{j}$ constrained up to $j$-th moment
Full Bayesian posterior $\Pi(\cdot\mid\mathcal{F}_{n})$ specified all moments
Conformal / rank calibration marginal coverage none
The hierarchy is strict: specifying $j$ moments determines $k$-step predictives for $k\leq j$ (Corollary 4.2) but not for $k>j$ (Theorem 6.3 at order $j+1$). The middle tier corresponds to linear Bayes / prevision updating (cf. Goldstein (1975a)).
Remark 10.1 (Conformal row).
Conformal prediction achieves marginal coverage $\mathbb{P}(Y_{n+1}\in\hat{C}_{n})\geq 1-\alpha$ under exchangeability without specifying any moments. The resulting prediction sets are finite-sample valid and distribution-free, but set-valued; marginal coverage does not imply correct specification of predictive probabilities. This is a qualitatively different guarantee from predictive completeness and the two are not directly comparable.
For a predictive characterization of exchangeable laws via mixing measures, see Fortini and Petrone (2017).
10.2 Closure theorem
Definition 10.2 (Predictive completeness).
A martingale posterior is predictively complete if it determines $\mathbb{P}(X_{n+1:n+k}=x_{1:k}\mid\mathcal{F}_{n})$ for all $k\geq 1$ and all cylinder events.
Theorem 10.3 (Closure).
Let $(\theta_{n})$ be a $[0,1]$-valued martingale satisfying Assumption 3.4, with terminal value $\theta_{\infty}$. The following are equivalent:
  1. (i)
    The martingale posterior determines $\mathbb{E}[(1-\theta_{\infty})^{k}\mid\mathcal{F}_{n}]$ for all $k\geq 1$.
  2. (ii)
    The conditional law of $\theta_{\infty}$ given $\mathcal{F}_{n}$ is uniquely determined (a.s.).
  3. (iii)
    The conditional moment sequence $(\mathbb{E}[\theta_{\infty}^{j}\mid\mathcal{F}_{n}])_{j\geq 1}$ is uniquely determined (a.s.).
Proof.
[…]
A martingale posterior is predictively complete if and only if it uniquely determines the conditional law of $\theta_{\infty}$ given $\mathcal{F}_{n}$. Condition (1) is necessary but not sufficient: it fixes only $\mathbb{E}[\theta_{\infty}\mid\mathcal{F}_{n}]=\theta_{n}$ and leaves the conditional law underdetermined. On $[0,1]$, specifying the law is equivalent to specifying all moments (Hausdorff determinacy; Theorem A.1), which is equivalent to specifying all $k$-step predictives.
11 Optimal stopping and dynamic amplification
The moment hierarchy has direct consequences for sequential decision problems.
11.1 Setup
Fix $K\geq 2$. At time $n$, define the value of the multi-step prediction problem under the full posterior law:
$V_{n}\;:=\;\sup_{2\leq\tau\leq K}\;\mathbb{E}\!\left[(1-\theta_{\infty})^{\tau}\mid\mathcal{F}_{n}\right].$ (22)
The constraint $\tau\geq 2$ restricts attention to genuinely multi-step objectives; the case $\tau=1$ reduces to first-moment coherence already determined by the martingale condition. Under the mean-only plug-in, replace $\theta_{\infty}$ by its conditional mean $m_{n}=\mathbb{E}[\theta_{\infty}\mid\mathcal{F}_{n}]$:
$\widetilde{V}_{n}\;:=\;\sup_{2\leq\tau\leq K}\;(1-m_{n})^{\tau}.$ (23)
11.2 Strict value distortion
Proposition 11.1 (Exact value gap).
Let $\sigma_{n}^{2}:=\mathrm{Var}(\theta_{\infty}\mid\mathcal{F}_{n})$. If $\sigma_{n}^{2}>0$, then $V_{n}-\widetilde{V}_{n}=\sigma_{n}^{2}$.
Proof.
Since $x\mapsto(1-x)^{\tau}$ is decreasing on $[0,1]$ for each fixed $\tau\geq 1$, both suprema are attained at the lower boundary $\tau=2$. By the second-moment identity $\mathbb{E}[(1-\theta_{\infty})^{2}\mid\mathcal{F}_{n}]=(1-m_{n})^{2}+\sigma_{n}^{2}$ (equation (12)), we have $V_{n}=(1-m_{n})^{2}+\sigma_{n}^{2}$ and $\widetilde{V}_{n}=(1-m_{n})^{2}$. ∎
Corollary 11.2 (Stopping boundary distortion).
For $r\in(0,1)$, define the effective prediction horizon under the full posterior as $\tau^{*}_{n}(r):=\max\{2\leq\tau\leq K:\mathbb{E}[(1-\theta_{\infty})^{\tau}\mid\mathcal{F}_{n}]\geq r\}$, and the plug-in horizon as $\widetilde{\tau}^{*}_{n}(r):=\max\{2\leq\tau\leq K:(1-m_{n})^{\tau}\geq r\}$. There exist states $\mathcal{F}_{n}$ and thresholds $r$ for which $\tau^{*}_{n}(r)>\widetilde{\tau}^{*}_{n}(r)$.
Proof.
By Proposition 11.1, for each $\tau_{0}\in\{2,\ldots,K-1\}$, $\mathbb{E}[(1-\theta_{\infty})^{\tau_{0}}\mid\mathcal{F}_{n}]>(1-m_{n})^{\tau_{0}}$ when $\sigma_{n}^{2}>0$. Choose $r$ strictly between these two values; then $\tau_{0}$ lies in the Bayes feasible set but not the plug-in feasible set. ∎
The distortion $V_{n}-\widetilde{V}_{n}=\sigma_{n}^{2}$ vanishes at rate $O(1/n)$ as the posterior concentrates (Proposition 9.3), so predictive completeness is necessary for optimal sequential decisions.
12 Discussion
Three notions of inferential validity appear in the literature and bear on the results developed here. Blackwell admissibility requires that no other rule uniformly dominates a given rule; by Theorem 7.2, admissible rules are generalized Bayes. Anytime-valid inference controls type-I error at arbitrary stopping times via e-processes, nonnegative supermartingales under $H_{0}$ (Ville, 1939; Grünwald et al., 2024). Conformal validity achieves marginal coverage $\mathbb{P}(Y_{n+1}\in\hat{C}_{n})\geq 1-\alpha$ under exchangeability, finite-sample and distribution-free but set-valued. These notions address distinct inferential objectives, and multi-step predictives—polynomial functionals of $\Pi(\cdot\mid\mathcal{F}_{n})$—fall under the scope of Theorems 6.3 and 10.3.
For one-step prediction ($k=1$), first-moment coherence suffices: $\mathbb{P}(X_{n+1}=1\mid\mathcal{F}_{n})=\theta_{n}$.
For $k\geq 2$, the martingale condition alone does not uniquely identify $k$-step predictives (Theorem 6.3). A martingale posterior construction that specifies the full conditional law of $\theta_{\infty}\mid\mathcal{F}_{n}$ is predictively complete (Theorem 10.3); one that specifies only the update rule $\theta_{n}\mapsto\theta_{n+1}$ leaves multi-step predictives not uniquely determined, and the plug-in $(1-\theta_{n})^{k}$ is inadmissible for $k\geq 2$ (Corollary 7.4).
Many practical martingale-posterior constructions do in fact specify a full conditional law for $\theta_{\infty}\mid\mathcal{F}_{n}$, in which case the closure condition of Theorem 10.3 is automatically satisfied.
Example 12.1 (Diagnostic application).
Consider a martingale posterior defined by the update $\theta_{n+1}=\theta_{n}+\gamma_{n}(X_{n+1}-\theta_{n})$ for a learning rate sequence $\gamma_{n}\to 0$. This satisfies (1) by construction. If $\gamma_{n}=1/(n+c)$ for some $c>0$, then $\theta_{n}=(S_{n}+c\theta_{0})/(n+c)$, which is the posterior mean under a $\mathrm{Beta}(c\theta_{0},c(1-\theta_{0}))$ prior. The conditional law of $\theta_{\infty}$ is then $\mathrm{Beta}(c\theta_{0}+S_{n},\,c(1-\theta_{0})+n-S_{n})$: fully specified, and all predictives are determined.
For other choices of $\gamma_{n}$ (such as $\gamma_{n}=n^{-\alpha}$ with $\alpha\neq 1$), the limiting conditional law is typically not a Beta distribution and may not be analytically available. The martingale condition still holds, but without the conditional law, the practitioner cannot compute $\mathbb{E}[(1-\theta_{\infty})^{k}\mid\mathcal{F}_{n}]$ for $k\geq 2$. In such cases, bootstrap or resampling approximations to the conditional law are needed to recover multi-step predictive coherence.
In classical Bayesian inference, the likelihood $\theta^{S_{n}}(1-\theta)^{n-S_{n}}$ updates $\Pi$ to $\Pi(\cdot\mid\mathcal{F}_{n})$, determining all moments. Goldstein (1975b) made this mechanism explicit: when the posterior mean is linear in the data, the likelihood moments $(m_{i})$ bootstrap from the first prior moment to all higher ones via a recursive identity. That recursion is the structural ingredient the martingale condition omits. Equivalently, the likelihood/prior specification supplies the typical-set geometry (KL/entropy rates) for learning along a single path (Polson and Zantedeschi, 2025). Absent this specification (whether via likelihood or alternative construction), multi-step predictives are not uniquely determined.
The moment hierarchy rests on the classical theory of moment problems (Hausdorff (1921); Feller (1971)). Hausdorff determinacy (that a measure on $[0,1]$ is uniquely determined by its moment sequence) is the key structural tool, used both to establish injectivity of the $k$-step run probabilities (Theorem 4.3) and to prove the closure theorem (Theorem 10.3). Compact support is essential: the moment problem is generically indeterminate for unbounded models (Remark 4.5), where analogous results remain open.
Fong, Holmes, and Walker (2023) introduced the martingale posterior as a general-purpose alternative to likelihood-based updating. Their framework specifies a predictive resampling scheme that, in implementations, often amounts to specifying a full conditional law for $\theta_{\infty}\mid\mathcal{F}_{n}$. The insufficiency results of Section 6 apply to the abstract martingale condition, not to specific implementations that go beyond it. The closure theorem (Theorem 10.3) and the optimal stopping distortion (Section 11) together characterize the full consequences of first-moment versus full-law specification.
Goldstein and Wooff (1998) provide a linear-Bayes treatment of adjusting exchangeable beliefs; their framework deliberately stops short of full distributional specification. Theorem 10.3 identifies the predictive consequences of that choice. The related contributions of Polson and Zantedeschi (2025) and Datta and Polson (2025) to finite-moment prediction and empirical Bayes are placed in context in the Introduction. Bayesian predictive inference via c.i.d. sequences, without a fully specified prior, is developed in Berti, Dreassi, Pratelli and Rigo (2021); Fortini and Petrone (2017) give a predictive characterization of exchangeable laws in a related setting. Leisen, Pratelli and Rigo (2025) establish that $X$ is c.i.d. if and only if $(\alpha_{n}(f))$ is a martingale for every bounded Borel $f$; the martingale posterior of Fong, Holmes and Walker (2023) constrains only $\alpha_{n}(\mathrm{id})=\theta_{n}$, and the present results quantify the predictive consequences of that gap through the moment hierarchy. In a complementary direction, Battiston and Cappello (2025) introduce almost conditional identically distributed (a.c.i.d.) sequences, in which predictive distributions are measure-valued almost supermartingales parametrized by a sequence of deviation parameters; where the present results identify what the strict martingale condition underdetermines, their framework broadens the class of admissible predictive schemes beyond the c.i.d. setting.
Several open questions emerge from this work.
Minimal sufficient conditions. The closure theorem (Theorem 10.3) states that predictive completeness requires the full conditional law. A natural question is whether there exist finitely many moment constraints (beyond the first) that yield “approximate” coherence in a precise quantitative sense. The second-order bound (Proposition 6.4) implies that specifying $\mu_{1}$ and $\mu_{2}$ controls $k$-step predictives up to third-order remainders, but the precise minimax rates over the class of posteriors with given $(\mu_{1},\mu_{2})$ are unknown.
Beyond Bernoulli. The moment hierarchy exploits the compact support $[0,1]$ via the Hausdorff moment theorem. For exponential families with non-compact natural parameter space (Gaussian mean, Poisson rate), the moment problem is generically indeterminate (Remark 4.5). Whether analogous insufficiency results hold, and what replaces the Hausdorff determinacy, is open. Carleman’s condition provides a sufficient criterion in the one-dimensional case; the multivariate setting requires additional structure.
Computational implications. When predictive completeness is needed, one must specify the conditional law of $\theta_{\infty}\mid\mathcal{F}_{n}$. Whether approximate methods (variational, ABC) can be disciplined by moment constraints to achieve practical completeness is an open question.
Connection to calibration. Conformal prediction achieves marginal coverage without moment constraints (Table 4). Whether there exists a unified framework that interpolates between marginal calibration and full predictive coherence, parameterised by the number of constrained moments, would connect the conformal and Bayesian literatures.
In summary, the martingale posterior framework provides a flexible mean-calibrated updating mechanism; the present results characterize the additional structure required for predictive completeness. Martingale coherence (1) constrains only the first conditional moment. Predictive completeness requires the full conditional law, equivalently all moments on $[0,1]$ (Theorem 10.3). The Bayes posterior supplies this; the martingale condition alone does not.
Appendix A Hausdorff moment theorem
Theorem A.1 (Hausdorff).
A sequence $(\mu_{k})_{k\geq 0}$ with $\mu_{0}=1$ is the moment sequence of a probability measure on $[0,1]$ if and only if it is completely monotone: $(-1)^{m}\Delta^{m}\mu_{k}\geq 0$ for all $k,m\geq 0$, where $\Delta$ is the forward difference operator. The moment sequence uniquely determines the measure.
See Hausdorff (1921) and Feller (1971), Chapter VII.
Appendix B General functional non-identification
Theorem B.1 (General nonlinearity).
Let $f:[0,1]\to\mathbb{R}$ be any non-affine continuous function. For any $m\in(0,1)$, there exist probability measures $\nu_{1},\nu_{2}$ on $[0,1]$ with $\int\theta\,\mathrm{d}\nu_{1}=\int\theta\,\mathrm{d}\nu_{2}=m$ but $\int f\,\mathrm{d}\nu_{1}\neq\int f\,\mathrm{d}\nu_{2}$. Therefore $\mathbb{E}[f(\theta)\mid\mathcal{F}_{n}]$ is not determined by $\mathbb{E}[\theta\mid\mathcal{F}_{n}]$.
Proof.
Since $f$ is non-affine and continuous, there exist $a<b$ in $[0,1]$ such that $f$ is not affine on $[a,b]$. For $m\in(a,b)$, set $\nu_{1}=\delta_{m}$ and $\nu_{2}=\lambda\delta_{a}+(1-\lambda)\delta_{b}$ with $\lambda=(b-m)/(b-a)$. Both have mean $m$. By strict Jensen, $\int f\,\mathrm{d}\nu_{2}\neq f(m)=\int f\,\mathrm{d}\nu_{1}$. For $m\notin(a,b)$, a three-point mixture with positive mass on $[a,b]$ achieves the same conclusion by continuity and non-affinity of $f$. ∎
The case $f(\theta)=(1-\theta)^{k}$ recovers Lemma 6.1. The case $f(\theta)=\mathbf{1}[\theta\leq t]$ shows that even the posterior CDF at a single point is not determined by the mean; the case $f(\theta)=-\theta\log\theta-(1-\theta)\log(1-\theta)$ (Bernoulli entropy, strictly concave) gives the same conclusion for expected entropy.